An indicator is a weak acid and pH range of its colour is 3.1 to 4.5. If the neutral point of the indicator lies in the centre of the hydrogen ion concentration corresponding to the given pH range, calculate the ionisation constant of the indicator.

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
HA +H2O-------------H3O+ + A-
K=(A-)(H+)/(A-)
at equilibrium concentration of HA and A- will be equal
K=(H+)
(H+)=-log(3.8)=1.58*10^-4
k=1.58*10^-4